1. How much would you need to deposit in an account now in order to have $5000 in the account in 15 years? Assume the account earns 5% interest compounded monthly.

2. You currently have $7,700 (Present Value) in an account that has an interest rate of 3% per year compounded quarterly (4 times per year). You want to withdraw all your money when it reaches $18,480 (Future Value). In how many years will you be able to withdraw all your money?

3. Suppose you want to have $500,000 for retirement in 35 years. Your account earns 10% interest. How much would you need to deposit in the account each month?
Answer #1

Using financial calculator

1.
N=5*12=60
I/Y=5%/12
FV=-5000
PMT=0
CPT PV=3896.02695158484

2.
PV=-7700
FV=18480
I/Y=3%/4
PMT=0
CPT N=117.16635422342

Years=117.16635422342/4=29.29159

3.
FV=-35000
I/Y=10%/12
PV=0
N=12*35
CPT PMT=9.21868229799008